### v3

# Towards Clean-Label Backdoor Attacks in the Physical World

Abstract
Deep Neural Networks (DNNs) are shown to be vulnerable to backdoor poisoning attacks, with most research focusing on digital triggers – special patterns added to test-time inputs to induce targeted misclassification. Physical triggers, natural objects within a physical scene, have emerged as a desirable alternative since they enable real-time backdoor activations without digital manipulation. However, current physical backdoor attacks are dirty-label, meaning that the injected poisoned inputs have incorrect labels, making them susceptible to filtering by human inspection. Therefore, in this paper, we explore a new paradigm of backdoor attacks, Clean-Label Physical Backdoor Attacks (CLPBA), via experiments on two collected datasets of facial recognition and animal classification. Our study reveals that CLPBA could be a serious threat with the right poisoning algorithm and physical trigger. Notably, unlike digital backdoor attacks that exploit memorization to plant backdoors in models, CLPBA works by embedding the feature of the trigger distribution (i.e., the distribution of trigger samples) into the poisoned images through the perturbations. We also find that representative filtering and firewall defenses are not effective against CLPBA since CLPBA fundamentally breaks the core assumptions behind these defenses.
1 Introduction
The development of DNNs has led to breakthroughs in various domains, such as computer vision, natural language processing, speech recognition, and recommendation systems [5, 9, 34, 26]. However, training large neural networks requires a huge amount of training data, encouraging practitioners to use third-party datasets, crawl datasets from the Internet, or outsource data collection [16, 43]. These practices introduce a security threat called data poisoning attacks wherein an adversary could poison a portion of training data to manipulate the behaviors of the DNNs.
One line of research in data poisoning is backdoor attacks, in which the attackers aim to create an artificial association between a trigger and a target class such that the presence of such trigger in samples from the source class, or multiple source classes, will cause the model to misclassify as the target class. The backdoored model (i.e., the model trained on poisoned samples) behaves normally with ordinary inputs while misclassifying trigger instances (i.e., instances injected with the trigger), making backdoor detection challenging. For example, Gu et al. [16] show that a backdoored traffic sign classifier has high accuracy on normal inputs but misclassifies a stop traffic sign as “speed limit” when there is a yellow square pattern on it.
Most backdoor attacks employ digital triggers, special pixel patterns digitally added at inference time to cause misclassification. Another less explored line of research investigated the use of natural objects within a scene as physical triggers. Physical triggers have two major advantages over digital triggers: they appear more natural when injected into the inputs and they do not require digital manipulation during inference. This makes backdoor attacks on real-time tasks such as facial recognition and object detection possible. However, existing physical backdoor attacks are dirty-label, meaning that the labels of trigger samples (i.e., images containing the trigger) are converted to the target-class label. The inconsistency between image contents and labels allows data inspectors to easily remove trigger samples from the dataset, especially when such samples contain a visible trigger (e.g., sunglasses) as a distinctive feature. Thus, this paper addresses the research question: ”Is it feasible to execute a successful physical backdoor attack by injecting a small number of poisoned instances that are clean-label and triggerless into the training set?
Figure 1: General pipeline of CLPBA. With access to the training dataset and trigger samples from the source class, the attacker uses the attacker model to optimize perturbations that are subsequently added to a small number of target class samples. At inference time, the victim model trained on these perturbed samples will incorrectly classify the source-class samples with the trigger as the target class.
This paper introduces Clean-Label Physical Backdoor Attacks (CLPBA), characterized by:
  • Clean-label: The injected poisoned samples preserve the ground-truth labels instead of having the target labels.
  • Hidden-trigger: Unlike standard backdoor attacks, the poisoned samples of CLPBA do not explicitly contain the trigger but are perturbed with constrained noise. Thus, they are harder to detect since there is no visible difference between poisoned and clean samples.
  • Real-time: These attacks do not require digital manipulation at test time, allowing for applications in real-time environments, such as facial recognition systems or traffic sign detection for autonomous vehicles.
Since there are currently no public datasets to study physical backdoor attacks, we collected a large-scale facial classification dataset that contains 21,238 images of 10 volunteers and 7 triggers representing common facial accessories, following the design of [52]. We also collected 2000 cat images (500 clean images and 1500 trigger images) containing 3 physical triggers (tennis ball, phone, book) and then combined these images with an existing animal classification dataset [1]. We then conducted extensive experiments on these two datasets to study CLPBA. Our key contributions can be summarized as follows:
  1. 1.
    We introduce CLPBA, a new threat of physical backdoor attacks that optimize perturbations on clean-label samples to embed the backdoor, instead of injecting dirty-label trigger samples. Thus, the attacks can be both stealthy (clean-label, hidden-trigger) and applicable in real-time via physical triggers.
  2. 2.
    Our study reveals that, different from digital backdoor attack,s which exploit memorization to plant backdoors in deep nets, CLPBA attacks the model by embedding the features of the source-class trigger distribution (i.e., the distribution of trigger images from the source class that are to be misclassified) to the poisoned images.
  3. 3.
    We identify a trade-off between perceptual stealthiness and the effectiveness of CLPBA. To address this issue, we propose a new pixel regularization consisting of a soft version of $\ell_{\infty}$ constraint and Upwind [6] term to smoothen out the perturbations. Empirically, our regularization improves the convergence rate of poisoning algorithms, thereby enhancing the effectiveness and stealthiness of the poisoned samples.
2 Related Works
In Backdoor Attacks, the attacker uses a predefined trigger to insert a backdoor into a victim model by poisoning a small proportion of training data. Training on the poisoned data causes the model to misclassify instances containing the trigger as the attacker’s chosen target label. Generally, backdoor attacks can be divided into dirty-label and clean-label backdoor attacks.
Dirty-label attacks. The attacker enforces a connection between the backdoor trigger and the target class by adding the trigger to the training data and flipping their labels to the target class. Types of Dirty-label Backdoor Attacks include trigger-pattern attacks [16, 2, 8], sample-specific attacks [36, 28], and imperceptible attacks [28, 11, 12, 35, 31]. While dirty-label attacks achieve impressive performance with a small poisoning ratio, mislabelled poisoned samples are vulnerable to human inspection as their image contents visibly differ from target-class instances.
Clean-label backdoor attacks. A more stealthy approach involves directly poisoning target-class instances without label manipulation. The concept of clean-label backdoor attacks was pioneered by Turner et al. [50], who proposed using adversarial perturbations and GAN-based interpolation to obscure the natural, salient features of the target class before embedding the trigger. By effectively concealing the latent features with the perturbations, the model becomes reliant on the introduced trigger for classifying instances of the target class. The following works on Clean-label attacks can be divided into hidden-trigger and trigger-design attacks. In hidden-trigger attacks [41, 46], the trigger is hidden from the training data and only added to test-time inputs of the source class to achieve the targeted misclassification. In trigger-design attacks [57, 23], the attackers aim to optimize trigger pattern(s) that represent the most robust, representative feature of the target class.
Physical backdoor attacks. Both clean-label and dirty-label backdoor attacks require digitally modifying inputs to insert the trigger. In practical scenarios, however, digital modification is not always possible, especially for real-time applications such as facial recognition and object detection. As an alternative, some works focus on using physical objects as triggers. Notably, Gu et al. [16] demonstrated that a yellow post-it note patched on a stop sign could cause the traffic sign to be misclassified as a speed limit. Similarly, Chen et al. [8] demonstrated an attack that fools a facial recognition system by blending the digital image of sunglasses to the training data and using the physical object (of the same sunglasses) to fool the model at inference. Wenger et al. [52] conducted an empirical study on the effectiveness of physical objects as backdoor triggers by collecting a dataset with 3,205 images of 9 facial accessories as potential backdoor triggers. These studies on physical backdoor attacks are limited to dirty-label techniques, which is not the focus of our research.
Of particular relevance to our work, Zeng et al. [57] proposed a clean-label backdoor attack that can work in a physical context. However, the paper lacks detailed information on experimental design, dataset, and experiments, which hinders definitive conclusions about the attack’s performance in real-world scenarios. Moreover, their approach differs from CLPBA because it designs a conspicuous trigger pattern instead of exploiting a natural object. Thus, we are the first to comprehensively study the performance of clean-label backdoor attacks with physical triggers on two tasks, facial recognition and animal classification.
3 Natural Physical Backdoors
For a backdoored DNN, the backdoor trigger can be understood as the strong feature for classification of the target class [38, 25, 49]. In real-world datasets, the model could “overfit” to a specific feature in the target-class data. This class-wise feature could be natural textures (e.g., eye, nose, mouth of a human), or an artificial object injected into the scene, such as a special pair of eyeglasses or a face mask. In this section, we explore the question: “Could a dataset contain a natural backdoor that the attacker can exploit?
To answer this question, we handpick a target class T (class 6 in our dataset), whose dataset contains black eyeglasses in the majority of the images, and a source class S (class 0 in our dataset) that does not wear any accessories on the face. We gradually inject 0, 12, and 23 target images with trigger sunglasses (0%, 5%, and 10% of the target-class dataset) into the training dataset.
Figure 2: ASR (%) on ResNet50 and VGG16 models when the attacker injects target-class images with sunglasses into the dataset.
For every injection attempt, we record the attack success rate (ASR) on a held-out dataset containing 121 images of the S wearing sunglasses. Interestingly, even with 0 injected images, the ASR could reach 26.21 % for ResNet50 and 15.51% for VGG16 (Figure 2). Injecting only 12 images (5% of the target class training dataset) increases the ASR up to 75.21% for ResNet50 and 54.7% for VGG16.
[…]
$\text{dist}(x,\widetilde{\mathcal{D}}_{s})=1-\frac{x^{\top}\mu}{||x||\cdot||\mu||}$ (1)
where $\mu=\frac{1}{|\widetilde{D}_{s}|}\sum_{x\in\widetilde{D}_{s}}f(x)$ is the empircal mean of feature representations for $\widetilde{\mathcal{D}}_{s}$.
Table 1: ASR and ACC of ResNet50 before and after filtering.
ASR (%) ACC (%)
Before filtering 61.54 99.85
After filtering 5.98 99.68
Random filtering 69.96 99.53
Results from Table 1 show that training on filtered data significantly lowers ASR from 61.54% to 5.98%, which is lower than when no trigger instances are injected ($26.21\%$). As a comparison, randomly filtering the same number of samples in the target class, and the ASR increases to $69.96\%$, possibly because clean samples are filtered.
Figure 3: From left to right, the box plots of cosine distances (Equation 1) in the feature space of samples from $\widetilde{\mathcal{D}}_{s}$ to $\widetilde{\mathcal{D}}_{t}$, $\mathcal{D}_{t}$, and $\mathcal{D}_{s}$.
In Figure 3, we plot the feature distances of samples from 3 distributions to $\widetilde{\mathcal{D}}_{s}$. From the first two boxes, we can see that the existence of a physical trigger shifts the distance of samples in $\widetilde{\mathcal{D}}_{t}$ closer to $\widetilde{\mathcal{D}}_{s}$. The average distance of sample from $\widetilde{\mathcal{D}}_{t}$ to $\widetilde{\mathcal{D}}_{s}$ is even smaller than the average distance of sample from $\mathcal{D}_{s}$ to $\widetilde{\mathcal{D}}_{s}$. This means that when the model overfits to the sunglasses feature of the target class dataset, inference-time samples from $\widetilde{\mathcal{D}}_{s}$ will be mapped to a region closer to $\widetilde{\mathcal{D}}_{t}$ than $\mathcal{D}_{s}$ in the feature space (Figure 3), resulting in mis-classification.
(a) Clean samples (w/o trigger)
(b) Trigger samples
Figure 4: Grad-CAM [44] attention visualization for clean ResNet50 model on (a) clean samples and (b) trigger samples. The first row is source class 3; the second row is source class 4. The attention is targeted at ground truth labels (i.e, the first row is target 3 and the second row is target 4).
Table 2: GM attack performance for source classes 3 and 4 (target: class 6). Clean model loss is measured on inference-time trigger samples from the source class.
Source $\to$ Target Clean Loss ASR (%)
3 $\to$ 6 0.6855 99.0 (0.8)
4 $\to$ 6 0.0038 86.8 (2.1)
[…]
4 Clean-Label Physical Backdoor Attack
4.1 Problem Formulation
The effectiveness of the Naive attack, as introduced in Section 3, is highly dependent on the combination of source class, target class, and physical trigger and thus may not generalize to realistic settings. Furthermore, the existence of a physical trigger in inserted poisoned images may raise suspicion. In this section, we introduce the problem formulation and poisoning algorithms of CLPBA, which crafts triggerless poisoned samples that indirectly plant the backdoor (activated by a physical trigger) into the models.
Notations. Denote $y_{t}$ as the label of target class. Let $D=\bigcup^{k}_{i=1}D_{i}$ be the training dataset with $k$ classes. Let $\mathcal{D}_{i}$, $\widetilde{\mathcal{D}}_{i}$ be the distribution of instances in class $i$ without and with the physical trigger. Let $F\colon\mathcal{X}\rightarrow R^{k}$ be the classifier model, with parameters $\theta$, that gives prediction scores (i.e., logits) for each of $k$ classes given an input $x\in\mathcal{X}$. $F$ can be decomposed as $F=f\circ g$ with $f$ as the feature extractor and $g$ be the classifier. The poison pair $s-t$ is the pair of source class and target class with index labels.
The attacker aims to inject a poison set ${P}$ to the training set ${D}$ such that the victim model trained on ${D}^{p}:={D}\cup{P}$ will misclassify trigger samples from the source class as the target class. To achieve this goal, the attacker needs to solve the following bilevel optimization problem:
$\underset{\delta\in C}{\text{min}}\;L_{adv}(\delta)=\mathbb{E}_{(\tilde{x}_{s}\sim\widetilde{\mathcal{D}}_{s})}\Bigl{[}\mathcal{L}\bigl{(}F(\tilde{x}_{s};\theta(\delta)),y_{t}\bigr{)}\Bigr{]}$ (2)
$\theta(\delta)\in\operatorname*{arg\,min}_{\theta}\;L_{train}(\theta)=\sum_{x,y\in D^{p}}\mathcal{L}\bigl{(}F(x;\theta),y\bigr{)}$ (3)
where $C$ is the constraint set of the perturbations: $C=\{\delta\;|\>||\delta||_{\infty}<\epsilon\}$. $\mathcal{L}$ is the objective function and $\widetilde{\mathcal{D}}_{s}$ is the distribution of trigger instances from the source class. The poison set $P$ is a small subset of $D_{t}$. Formally, ${P}=\{x_{t}^{(i)}+\delta^{(i)},y_{t}\}^{K}_{i=1}$ contains $K$ poisoned target-class samples. The value $\epsilon$ controls the magnitude of the perturbations and thus the stealthiness of poisoned samples. We define the poison ratio $\alpha$ as the percentage of poisoned samples in the target-class data. Note that this may differ from previous works on backdoor attacks which define poison ratio over the entire training set. With the above formulation, we introduce three poisoning algorithms for CLPBA based on existing clean-label data-poisoning methods:
  • Feature Matching (FM) [41]: This method crafts perturbations that shift poisoned samples closer to source trigger distribution $\widetilde{\mathcal{D}}_{s}$ in the feature space, thus representing them as dirty-label samples with trigger:
    $\mathcal{A}_{FM}=\min_{\delta\in C}\|f(x_{t}+\delta)-f(\tilde{x}_{s})\|_{2}^{2}\quad$ (4)
    $\text{where }x_{t}\in P,\tilde{x}_{s}\sim\widetilde{\mathcal{D}}_{s}$ (5)
  • Meta Poison (MP) [21]: This method is based on meta-learning [13], where it approximates the inner objective $\theta(\delta)$ by unrolling $K$ (default to 2) SGD steps before computing outer objective:
    $\theta_{1}$ $=\theta_{0}-\alpha\nabla_{\theta}L_{train}(\theta_{0}\;|\;\delta_{i})$ (6)
    $\theta_{2}$ $=\theta_{1}-\alpha\nabla_{\theta}L_{train}(\theta_{1}\;|\;\delta_{i})$ (7)
    $\delta_{i+1}$ $=\delta_{i}-\beta\nabla_{\delta}L_{adv}(\delta_{i}\;|\;\theta_{2})\text{ s.t }\delta\in C$ (8)
  • Gradient Matching (GM) [46]: This method crafts perturbations that align the gradient of the model on the poisoned dataset $P$ with the gradient to optimize $L_{adv}$:
    $\mathcal{A}_{GM}$ $=\min_{\delta\in C}\left(1-\frac{\nabla_{\theta}L_{poi}\cdot\nabla_{\theta}L_{adv}}{\|\nabla_{\theta}L_{poi}\|\cdot\|\nabla_{\theta}L_{adv}\|}\right)$ (9)
    where $\text{where }L_{poi}$ $=\frac{1}{K}\sum_{i=1}^{K}L(F(x_{t}^{(i)}+\delta^{(i)};\theta),y_{t})$ (10)
4.2 Threat Model
Attacker’s capability. In our threat model, the training dataset is collected from multiple sources, and there exists an attacker (e.g., data contributors, data collectors) who can access and perturb a small portion of the data to manipulate the behavior of the victim model. Following previous works [41, 46], we assume a gray-box scenario where the attacker knows the architecture of the victim model but does not have access to its training process. The attack is one-to-one: any instances from the source class are misclassified as belonging to the target class with the presence of the trigger. For example, the source class is an employee in a company who can wear a special pair of sunglasses to fool a facial recognition system into classifying him as the CEO so that he could get an illegal privilege, such as access to confidential documents or high-ranking facilities. Due to the nature of the physical trigger, we argue that all-to-one attacks have low utility as this scenario could raise suspicion among model users; for example, benign participants wearing the trigger object could accidentally activate the backdoor behavior and report the issue to authority.
Figure 5: Physical triggers in the collected datasets
We further assume the attacker can select a trigger from a limited pool of physical objects and has access to trigger samples from the source class. This assumption is reasonable, as the source class could be the attacker or collaborators. In some cases, physical triggers, being natural objects, may already exist in the training dataset of the target class [53]. In this paper, we consider a realistic scenario where the target class dataset lacks natural triggers, and the attacker cannot directly inject trigger instances into it.
4.3 Data Collection
Our facial dataset contains 21,238 images of 10 volunteers (9 males, 1 female, all Asian) and 7 triggers. It consists of 3,344 clean images, 9,790 trigger images of 7 physical triggers, and 8,104 additional images to evaluate accidental backdoor activations (see Section 8). Our animal dataset is created by combining a public Kaggle dataset [1], containing 45 classes of mammals, with a new cat class containing 500 cat images. For this dataset, we collected 1,500 cat images with 3 physical triggers (tennis ball, phone, book) to study CLPBA. The images are captured in various indoor and outdoor settings, under varying weather conditions, and with diverse shooting angles and distances. Details of our dataset can be found in the Appendix B.
Table 3: ASR (%) of poisoning algorithms on selected source-target pairs, averaged over 10 runs with standard deviation in brackets.
Scenario Poison ratio $\alpha$ Naive MP FM GM
$\epsilon=8/225$ $\epsilon=16/255$ $\epsilon=8/225$ $\epsilon=16/255$ $\epsilon=8/225$ $\epsilon=16/255$
(a) Facial Recognition (sunglasses) trigger
finetuning 5% 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 0.6 (0.5) 5.8 (3.2) 22.7 (10.9) 95.6 (1.8)
10% 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 3.4 (0.5) 30.4 (1.5) 51.4 (17.8) 97.8 (0.2)
transfer 5% 0.0 (0.0) 2.2 (0.5) 1.8 (1.3) 81.8 (5.6) 77.8 (9.6) 91.2 (7.6) 92.5 (8.5)
10% 0.0 (0.0) 2.4 (1.8) 3.6 (1.2) 79.3 (2.3) 64.5 (25.7) 93.6 (6.9) 97.2 (2.8)
(b) Animal Classification (tennis) trigger
fine-tuning 5% N/A 0.0 (0.0) 0.0 (0.0) 0.3 (0.5) 0.2 (0.3) 8.2 (4.4) 9.6 (4.1)
10% N/A 0.0 (0.0) 0.0 (0.0) 1.0 (0.8) 0.2 (0.4) 20.5 (8.9) 17.8 (7.7)
transfer 5% N/A 1.1 (0.7) 3.2 (2.5) 34.5 (1.9) 29.7 (3.0) 62.0 (2.1) 75.6 (5.5)
10% N/A 1.5 (1.2) 2.8 (1.2) 33.6 (2.8) 34.3 (2.3) 92.8 (1.2) 88.4 (3.7)
5 Experiments
5.1 Experiment Settings
We split our datasets with a ratio of 8:2 for train and test data. We use 50% of source-class trigger images to craft poisoned samples, and the remaining 50% is used to evaluate ASR. Note that the prediction accuracy (ACC) on clean samples consistently exceeds 99% for facial recognition and 93% for animal classification across different attacks, so it is not reported in the paper. We choose ResNet50 [18], pre-trained on the VGGFace2 dataset [4], to be the victim model for Facial Classification, while ResNet18 pre-trained on Imagenet [40] is used For Animal classification. We evaluate the attacks under transfer learning and fine-tuning settings, two common settings in clean-label data poisoning [41, 46]. For transfer learning, the feature extractor is frozen and only the last linear layer is trained, while for fine-tuning, we set a low learning rate of 0.001 to train the victim model on the poisoned dataset. To enable a fair comparison across different poisoning algorithms, we standardize the poisoning hyperparameters, including the number of iterations, optimizer, batch size, learning rate, and other relevant parameters. For the Facial Classification task, we select the poison-pair $9-5$ (as class labels in our dataset) for our experiment. This poison pair is manually selected to avoid dataset bias, as explained in Section 3. For the Animal Classification task, we select the poison-pair cat-fox.
5.2 Baseline evaluation
From the experiment results demonstrated in Table 3, we have the following observations:
  • GM is the most effective poisoning algorithm. GM is especially effective for the fine-tuning scenario, which is a harder scenario compared to transfer learning since the weights of all layers of the victim model are updated. Matching the gradient update of $L_{adv}$ allows the GM perturbations to utilize the backward information of all layers, instead of only the last linear layer as in FM. MP is not effective because the poisoned samples overfit the updated parameters of the unrolled surrogate model in the inner loop, which greatly differ from the victim model.
  • Impact of $\epsilon$ and $\alpha$. While increasing either $\epsilon$ or $\alpha$ would improve the performance of attacks, we observe that $\epsilon$ plays a more important role than $\alpha$, especially for fine-tuning scenario. A high value of $\epsilon$ is critical for the poisoning objective to converge, while a high value of $\alpha$ does not guarantee that every poisoned sample is effective.
  • Effectiveness of CLPBA varies across datasets. As shown in Table 3, CLPBA is significantly less effective in Animal Classification than in Facial Recognition. This reduced performance is due to the tennis trigger being as dominant as the sunglasses trigger while also appearing in varying locations across the trigger samples, making it harder for the model to learn.
5.3 Analysis of CLPBA
Two unique attributes of CLPBA are that it is hidden-trigger and source-specific. The first attribute means the trigger is not present in the poisoned samples, while the second ensures only source-class samples can cause the model to misclassify with the trigger. CLPBA achieves these attributes by embedding features of the source-class trigger distribution, $\widetilde{\mathcal{D}}_{s}$, into the perturbations in the poisoned target samples. Specifically, the embedded features consist of adversarial features and trigger features. We define adversarial features as the non-robust features [24] that shift the representations of target instances toward the source-class distribution in the feature space, and trigger features as the high-level features of the physical trigger extracted by the neural network. The existence of adversarial features ensures the source-specific attribute, while reconstructed trigger features allow the model to indirectly learn the physical trigger as a feature of the target class.
Figure 6: Experiment to show the existence of adversarial features. We set a high $\epsilon$ of 32, $\alpha=2$, poison pair is $9-5$.
Existence of adversarial features. To justify the existence of adversarial features, we take inspiration from [24] to test whether the model trained on poisoned target samples could successfully predict source-class samples. We design the experiment as illustrated in Figure 6. Then, we filter out all source-class samples before relabelling all GM-poisoned samples to the source class. We then train the victim model from scratch on this modified dataset before evaluating the recall (true positive rate) of predicting source-class samples that have been filtered previously. To avoid randomness, we set a prediction threshold of 85% after the softmax function. Notably, the model trained on this modified dataset has a prediction accuracy of 75% on source-class samples. This shows that the perturbations on poisoned target instances contain adversarial features representative of the source-class distribution. To make sure that the trained model is not biased towards predicting out-of-distribution samples as the source class, we also evaluate the false positive rate of classifying 200 random face images from the CelebA dataset [32] as the source class, and the accuracy is 0%. The setup of this experiment can be found in Appendix D.
Existence of trigger features. By scaling the perturbations on GM-poisoned samples, we observe that the perturbations appear to reconstruct the shape of the physical trigger. Interestingly, setting a high value of $\epsilon$ even reconstructs the samples in $\widetilde{\mathcal{D}}_{s}$, as can be observed in Figure 8.
(a) Clean model
(b) Poisoned model
Figure 7: Feature space visualization.
Feature space analysis. In Figure 7, we project source-class samples (green), target-class samples (blue), trigger source samples (yellow), and poisoned target samples (red) in 2-dimension feature space of both clean and poisoned models where the first axis is $\mu_{s}-\mu_{t}$ and the second axis is $\theta_{s}-\theta_{t}$, the normal vector to the decision boundary between the source and target class. In both feature spaces, we can see that some poisoned target instances are close to the samples of $\widetilde{\mathcal{D}}_{s}$. For the clean model, $\widetilde{\mathcal{D}}_{s}$ overlaps with ${\mathcal{D}}_{s}$ in the feature space. For the poisoned model, however, $\widetilde{\mathcal{D}}_{s}$ is shifted towards $\widetilde{\mathcal{D}}_{t}$ in the feature space. From this observation, we can say that since poisoned target instances contain representative features of $\widetilde{\mathcal{D}}_{s}$, especially the trigger feature which is not a part of ${\mathcal{D}}_{s}$, training the model on these poisoned instances cause $\widetilde{\mathcal{D}}_{s}$ to move closer to the classification region of the target class in the feature space.
Figure 8: From left to right: tennis, fake-beard, and sunglasses triggers. First row: a sample in $\widetilde{D}_{s}$. Second row: GM perturbation.
For filtering defenses, we evaluate Spectral Signature (SS) [49], Activation Clustering (AC) [3], SPECTRE [17], and SCAN [48]. Table 6 shows that only SS is effective against CLPBA. Although we can see that AC and SPECTRE have high elimination rates (62.16% and 72.97%), the attack still achieves 99.7 % ASR. This shows that the GM poisoning objective (Equation 9) focuses on a small subset of effective poison instances to match the gradient of adversarial loss and distill the features from the trigger distribution. This is supported in Figure 7 where only a subset of poisoned samples overlap with the trigger distribution. Detecting such effective instances is generally difficult since they correlate strongly with each other and they transform the model’s decision boundary of the target class. Table 6 also shows that firewall defenses are not effective for CLPBAs because they either miss trigger samples (low TPR) or incorrectly filter out too many benign samples (high FPR). This behavior is expected since these defenses are built on the memorization property of backdoor attacks, which differs from the generalization property of CLPBA.
6 Improving the Effectiveness of CLPBA
6.1 Pixel regularization
Empirically, higher $\epsilon$ lead to a higher ASR since the poisoning objective can converge to a smaller value. However, the perceptual quality of poisoned samples is compromised since the poisoned images have high-frequency artifacts that are clearly visible (Figure 8). The drawbacks of the $\ell_{\infty}$ regularization on perturbations are that it is non-differentiable, and the projection on $\ell_{\infty}$-norm ball is only executed at the end of an optimization step, making backdoor and poison stealthiness two separate tasks. As a consequence, the perceptual quality of poisoned images is compromised because (1) the pixel values of perturbation are often pushed to the boundary of the norm ball, (2) the noise can appear in unnecessary regions of the image, and (3) the perturbations contain high-frequency artifacts. To improve the visual quality and convergence of the poisoning objective, we propose to replace the current $\ell_{\infty}$ regularization with a Pixel regularization containing a soft pixel-wise $\ell_{\infty}$ loss and discretized Total Variation Upwind (TVU) [6]:
$\mathcal{R}_{pix}(\delta)$ $=||\delta^{+}-M\odot\epsilon||_{1}+\alpha TVU(\delta;\beta=1)$ (11)
$TVU(\delta;\beta)$ $=\sum_{p}^{\text{all pixels}}\biggl{(}\sum_{q\in\mathcal{N}(p)}\big{(}\delta^{(p)}-\delta^{(q)}\big{)}^{2}\biggr{)}^{\frac{\beta}{2}}$ (12)
where $\delta^{+}$ takes the absolute value of every element in $\delta$, $M$ is a mask of pixel-wise embedding ratio: $M_{i,j,k}=\frac{x_{i,j,k}}{\max_{k}{x_{i,j,k}}}$, assuming $x_{i,j,k}\in[0,255]$. $M$ encourages the perturbation to align with the underlying image content. In Equation 12, $\mathcal{N}(p)$ are neighbouring pixels of pixel $p$. The weighting factor $\alpha$ is set to 5. Both terms in Equation 11 improve the smoothness of perturbations, with $TVU(\delta)$ minimizing the difference between adjacent pixels.
Side-by-side comparison between standard $\ell_{\infty}$ regularization and our pixel regularization is given in Figure 9 and Figure 10. Our pixel regularization leads to a high ASR even for small values of $\epsilon=[4,8]$, while the ASR for standard $\ell_{\infty}$ regularization is below 20%. At the same time, our pixel regularization also has higher PSNR values for the poisoned images at every $\epsilon$, indicating better perceptual stealthiness. Table 4 demonstrates the improved effectiveness of pixel regularization across different poison pairs.
(a) Standard $\ell_{\infty}$ regularization
(b) Pixel Regularization
Figure 9: Visualization between standard $\ell_{\infty}$-regularization and our Pixel Regularization. Poisoned samples are crafted with GM attack ($\alpha=0.1$, $\epsilon=16/255$)
Table 4: ASR (%) of GM attack + tennis with two regularizers.
Regularizer Fox Kangaroo Mongoose Sea-lion Tapir
$\ell_{\infty}$ regu 9.6 (4.1) 8.4 (0.8) 5.3 (0.4) 13.4 (2.9) 7.8 (5.0)
pix. regu 30.9 (10.4) 12.5 (1.6) 4.4 (1.0) 24.2 (2.3) 26.6 (5.2)
(a) Standard $\ell_{\infty}$ regularization
(b) Pixel Regularization
Figure 10: Comparison between standard $\ell_{\infty}$-regularization and our Pixel Regularization on different values of $\epsilon$. Higher PSNR indicates better perceptual quality of poisoned images.
6.2 Ensembling and Retraining
Following [46], we also apply ensembling and retraining for CLPBA methods in the fine-tuning setting of the Animal Classification task. In our experiment, we apply an ensemble of 3 ResNet18 models; each of these models will retrain on the poisoned dataset after every 200 attack iterations. The attacker optimizes poisoned samples with GM for 1000 iterations. Table 5 shows that applying these techniques significantly improves the attack success rate.
Table 5: ASR (%) of GM ($\alpha=0.1,\epsilon=16/255$) on the Animal Classification task with tennis trigger.
Method ASR (%)
Baseline 17.8 (7.7)
Retraining 25.5 (4.2)
Ensemble+Retraining 29.7 (3.1)
7 Defending against CLPBA
We evaluate the performance of CLPBA against representative filtering defenses (filter poisoned samples during training), and firewall defenses (filter trigger samples during inference). For the filtering defenses, we evaluate Elimination Rate (ER), the percentage of poisoned samples that are correctly filtered, and Sacrifice Rate (SR), the percentage of clean samples that are incorrectly filtered. For firewall defenses, we evaluate the True Positive Rate (TPR), the percentage of trigger source-class samples that are correctly filtered, and the False Positive Rate (FPR), the percentage of non-trigger samples that are incorrectly filtered. For details on these defenses, we refer readers to Appendix F.
Table 6: Defense evaluation on GM attack ($\alpha=0.1$, $\epsilon=16/226$, poison pair $7-0$). ASR is 99.7% without defenses.
Filtering Defenses
Metrics AC[3] SS[49] SPECTRE[17] SCAN[48]
ER (%) 62.16 97.30 72.97 0
SR (%) 15.77 19.77 19.50 0
ASR (%) 99.7 0.0 99.7 99.7
Firewall Defenses
Metrics STRIP[14] CD[20] IBD-PSC[19] Frequency[56]
TPR (%) 0.0 79.6 0.0 0.0
FPR (%) 6.7 65.3 16.7 0.9
8 Accidental Backdoor Activations
For CLPBA to be stealthy, we argue that it needs to be source-specific and trigger-specific. That is, (1) only the source class can use the trigger to manipulate the victim model’s prediction at inference time; (2) the trigger should be discriminative so that the victim model does not incorrectly recognize similar-looking objects as triggers. We introduce two new metrics, False Positive Rate, and Suspicion Rate, to test accidental backdoor activations:
  • Suspicion rate (SR): The measurement of whether the victim model incorrectly misclassifies trigger samples from non-source classes as belonging to the target class.
  • False positive rate (FPR): The measurement of whether samples from the source class can fool the victim model with false triggers (i.e., objects that resemble the attacker’s physical trigger).
Poison pair Trigger Loss type ASR ACC FPR SR
9-5 Fake beard Original 98.7 99.9 50.8 12.5
With repel 98.7 99.9 49.7 3.9
7-3 Sunglasses Original 97.5 99.3 0.7 13.9
With repel 97.5 99.4 0.9 2.5
Table 7: FPR and SR for GM attack. The set-up of our dataset to evaluate SR and FPR is given in Appendix B.
Between the two metrics, SR is more important because it directly impacts attack stealthiness if a benign client is misclassified as the target class with the trigger or similar-looking items. The issue of false triggers, measured by FPR, is challenging to solve as the attacker cannot anticipate all possible false triggers in the real world. However, the attacker can partially mitigate the issue of false source classes, as measured by SR, by adding repel terms to the adversarial loss $L_{adv}$ in Equation 3:
$L_{adv}=N\cdot\overbrace{\sum_{x\in\widetilde{D}_{s}}\mathcal{L}\bigl{(}F_{\theta}(x),y_{t}\bigr{)}}^{\text{adversarial loss}}$
[…]
Here, $N$ is the total number of classes, and $\widetilde{D}_{i}$ is the dataset of trigger instances in class $i$. The repel terms ensure trigger instances from non-source classes are repelled from the target class classification, enhancing trigger specificity to the source class. As shown in Table 7, incorporating the repel term in the GM objective (Equation 9) significantly reduces SR while maintaining high ASR.
One limitation of using the repel terms is that the attacker must possess trigger samples from non-source classes. To circumvent this, the attacker can patch clean images with digital triggers, mapping them to appropriate facial areas.
9 Conclusion
Through extensive experiments, our research underscores the vulnerability of DNNs to clean-label physical backdoor attacks. We then show that CLPBA works by embedding adversarial features and trigger features in the poisoned instances. Additionally, we propose a novel pixel regularization to improve the effectiveness and stealthiness of poisoned samples. A significant limitation identified is accidental backdoor activations of the backdoored model due to false triggers or non-source classes, which could compromise attack stealthiness. To address this, we propose incorporating repel terms into the adversarial objective to enhance trigger specificity to the source class.
Supplementary Material
Appendix A Dataset
A.1 Ethics & Data Protection
IRB approval. Before conducting our study, we submitted a ”Human Subjects Research – Expedited Review Form” to our country’s Institutional Review Board (IRB). Our study received approval from the chairman of the institutional ethical review board under decisions No 24/2016/QD-VINMEC, No 23/2016/QD-VINMEC, and No 77/2021/QD-VINMEC. We prepared a consent form beforehand to ensure transparency in the procedure of dataset collection. All 10 volunteers in our dataset provided explicit written consent for us to collect the dataset and use the images for research purposes, including permission to use the captured images in the research paper.
Dataset protection. In adherence to strict ethical standards and privacy considerations related to the sensitive nature of the human face dataset, our research follows a comprehensive protocol to protect the privacy and confidentiality of the collected data. To safeguard the data, all images are securely stored on a protected server, with access restricted solely to the authors for research purposes. Additionally, all images in the paper are partially obscured to ensure that the identities of our volunteers are not exposed.
A.2 Dataset Collection
Due to the lack of publicly available datasets to study physical backdoor attacks, we collect a facial recognition dataset following the design consisting of approximately 21,238 images of 10 volunteers (9 males, 1 female, all Asian) and 7 triggers. In total, our dataset contains 3,344 clean images, 9,790 trigger images of 7 physical triggers. To test unwanted backdoor activations (false positive rate and suspicion rate), we collect 8,104 additional images of 10 objects to evaluate unwanted backdoor activation. To reflect real-world conditions, the dataset was constructed in 2 months so that the images could be captured in various indoor/outdoor settings, under varying weather conditions, and with diverse shooting angles and distances. All photos are RGB and of size (224,224,3), taken using a Samsung Galaxy A53 and Samsung Galaxy S21 FE.
Table 8: Our assessment of chosen physical triggers
Trigger Stealthiness Location Size
Yes No On-face Off-face Small Medium Big
Earrings
Fake Beard
Sticker
Facemask
Hat
Sunglasses
Headband
In our dataset, we choose triggers based on 3 criteria: Stealthiness (Does the trigger look natural on human face?), Location (Is the trigger on-face or off-face?), Size (How big is the trigger?). With these criteria, we select 7 triggers with evaluations given in Table 8.
A.3 Dataset Preprocessing
After collecting the images, we utilize a pre-trained MTCNN [58] (Multi-task Cascaded Convolutional Networks) model to detect and crop the face area. This preprocessing step ensures that the face is the focal point of each image, effectively removing any background noise. The cropped face regions are then resized to a standard dimension of $224\times 224$ pixels.
Figure 11: Visualization of the face detection process. Left: Original image. Right: Processed image with face area cropped and resized to $224\times 224$ pixels.
A.4 Dataset Structure
The train and test data are trigger-free images (the participants do not wear any trigger items) of 10 classes corresponding to 10 volunteers. The train and test data are split in a $80:20$ ratio.
False triggers. To evaluate the sensitivity of poisoned models to false triggers, as measured by false positive rate (FPR) (see Appendix B.4), we select three additional false triggers for each of the main triggers. Because it is hard to find a representative set of false triggers for each of the main triggers, we select the false triggers by changing the sizes, colors, and shapes of the objects (e.g, we change the shape of earrings and hats, change the colors of the headband, change the size of sticker). We also select objects that have similar locations on the face as false triggers (e.g., headbands are selected as false triggers for the hat trigger and the sunglasses trigger). In Figure 20, each row corresponds to a main trigger on the left, followed by three false triggers on the right.
Remark. It is important to note that the selection of false triggers in our work is entirely subjective. The purpose is to show that the poisoned models are sensitive to false triggers in a physical backdoor attack. Therefore, a higher FPR of a trigger does not necessarily mean that that trigger is less specific and stealthy than other triggers.
A.5 Animal Classification Dataset
Besides facial recognition, we also conduct an extended evaluation of CLPBA on animal classification. We collected 2000 cat images with 500 trigger-free images and 500 images for each of the triggers, which are tennis balls, mobile phones, and books. The trigger-free cat images are resized to $224\times 224$ and then concatenated to an existing animal classification dataset on Kaggle. Visualization of trigger images for this dataset is given in Figure 19  [^1] https://www.kaggle.com/datasets/asaniczka/mammals-image-classification-dataset-45-animals.
Appendix B Experiment Settings
B.1 Computational Resources.
All the experiments are conducted on the two servers. The first one with 7 RTX 3090 Ti 24 GB GPUS, and the second one has 6 RTX A5000 24 GB GPUS. The code is implemented in Pytorch. We develop the framework based on the previous works [46, 15]. We use Pytorch as the framework for our model training.
B.2 Training Settings.
We evaluate the attacks on four Convolutional Neural Networks for facial recognition: ResNet50 [18], VGG16 [45], MobileNetV4 [42], Inception-ResNet [47]. We choose fine-tuning as the training method for the victim model since it is a widely adopted training method for its speed of convergence and accuracy. Compared to transfer-learning, Fine-tuning and training from scratch are shown to be hard scenarios for clean-label attacks [41, 46, 15]. Our models are pre-trained on VGGFace2 [4], a large-scale facial dataset containing 3.31 million images divided into 9131 classes. Since BLC requires an adversarially trained model, we perform fast adversarial training [54] for the pre-trained model on our dataset before crafting poisons. We do not apply any data augmentations to the training pipeline. We only standardize the training data using z-score normalization with the mean and standard variation calculated from the data. We observe that the training models converge very quickly and achieve the validation accuracy of over $99\%$ and $93\%$ by epoch 10 for Facial Recognition and Animal Classification tasks.
Table 9: Training Hyperparameters
Hyperparameters Value
Optimizer Stochastic Gradient Descent (SGD) [39]
Learning rate (lr) $0.001$
Decay rate $45e-$ 4
Batch size $64$
Training epochs $40$
Learning rate scheduler Decrease lr by $90\%$ at epoch 14, 25, 36
Table 10: Hyperparameters for Poisoning Algorithms
Hyperparameters Value
Optimizer Signed Adam
Learning rate (lr) $0.1$
Attack iterations $250$
Learning rate scheduler
Cosine Annealing [33]; the learning
rate decreases to 0.0001 at the last epoch
Trigger opacity (BLC) $32/225$
Inner SGD steps (MP) $2$
B.3 Attack Settings.
Our experiments follow the gray-box scenario in which the attacker model has the same architecture and is trained on the same dataset as the victim model but has different weight initialization, and the attacker cannot access and modify the weights and training procedure of the victim model. The dataset of trigger samples is split 50-50 into train and test sets. The train set is used for crafting poisons while the test set is used to validate the attacks. Table 10 gives the hyperparameters for poisoning algorithms.
In the original Gradient Matching attack, Souri et al. [46] proposed retraining the attacker model and selecting poison samples with maximum gradients to increase ASR. For a fair comparison between poisoning algorithms, however, we do not include such factors in our settings. In our experiment, the attacker selects the poison samples randomly.
B.4 Evaluation Metrics.
To evaluate the performance of CLPBA, we adopt two standard metrics for backdoor attacks:
  • Attack success rate (ASR): The proportion of examples in a held-out test set of source class wearing the physical trigger that the model misclassifies as target class.
  • Accuracy (ACC): The model’s prediction accuracy on clean, ordinary test-time data when the model is poisoned.
Appendix C Additional Experiments
C.1 Blended Label-Consistent Attack
For completeness, in our evaluation, we also evaluate Blended Label-Consistent attack [50, 8] which works by perturbing the poisoned samples with adversarial noise to make the salient features of the samples harder to learn before injecting the digital trigger, forcing the model’s classification to depend on the trigger.
Following [50], we first train an adversarial-trained model, before using the poisoning method in [50] to craft adversarial examples. Then we insert the digital trigger $p$ (e.g., digital sunglasses) via a perspective transformation $T$ that maps the trigger to the appropriate facial area using facial landmarks. The trigger is inserted with an opacity $\tau$.
$x_{t}$ $\leftarrow x_{t}+\operatorname*{arg\,max}_{\|\delta\|_{\infty}\leq\epsilon}\mathcal{L}(F(x_{t},\theta),y_{t})$ (13)
$x_{t}$ $\leftarrow(1-\tau)\cdot x_{t}+\tau\cdot T(p)$ (14)
The procedure of BLC is illustrated in Figure 12. Compared to random placement and fixed center placement of the trigger, our adaptive trigger placement method with facial landmarks has a higher attack success rate as it preserves the spatial relationship between the trigger and the face.
Figure 12: Procedure of BLC with $\tau=1$. We first add adversarial perturbations to the poisoned instance. Then, using facial landmarks, we map the trigger to the appropriate facial area with perspective transformation.
C.2 Experiments on Different Models
Table 11: CLPBA ($\alpha=0.1$, $\epsilon=16/226$) on VGG16
Poison pair Naive BLC MP FM GM
1-7 63.8 (4.4) 2.3 (0.5) 0.3 (0.5) 82.3 (12.5) 100.0 (0.0)
3-6 98.6 (1.4) 95.7 (1.4) 39.6 (9.9) 100.0 (0.0) 100.0 (0.0)
5-1 42.0 (14.2) 0.3 (0.5) 0.9 (0.0) 25.9 (15.8) 86.7 (1.9)
6-8 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 1.4 (1.7) 0.0 (0.0)
7-3 4.0 (3.0) 0.0 (0.0) 0.0 (0.0) 89.1 (1.6) 85.2 (7.3)
9-5 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 6.3 (4.0) 6.3 (4.0)
ACC (%) of clean model 98.6 (0.3)
Table 12: CLPBA ($\alpha=0.1$, $\epsilon=16/226$) on Inception-ResNet
Poison pair Naive BLC MP FM GM
1-7 5.4 (7.9) 0.3 (0.5) 13.7 (9.9) 0.3 (0.5) 96.0 (3.2)
3-6 61.4 (4.7) 48.8 (15.6) 96.1 (2.2) 80.2 (1.7) 100.0 (0.0)
5-1 4.3 (1.9) 0.9 (0.0) 0.0 (0.0) 2.8 (1.9) 98.5 (1.9)
6-8 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 96.4 (0.5)
7-3 0.3 (0.5) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 97.8 (2.1)
9-5 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 88.9 (6.9)
ACC (%) of clean model 99.8 (0.1)
We further evaluate the performance of CLPBA on 6 random poison pairs, on VGG16 and Inception-ResNet model architecture for the Facial Recognition task in Table 11 and Table 12.
C.3 Scaling the Dataset Size
To test the effect of dataset size and the number of classes on the performance of CLPBA, we scale our dataset by merging with two benchmark datasets, PubFig [27] and CelebA [32], with the following configurations:
  • PubFig dataset: We merge our dataset with 10, 20, 30, and 40 classes of PubFig with the highest number of images. The average number of images per class in the Pubfig dataset is approximately 293, which is comparable to our dataset. However, it is important to note that many images in this dataset have been corrupted due to incorrect download links.
  • CelebA dataset: We merge our dataset with 90, 190, and 390 classes of CelebA with the highest number of images.
Table 13: Merged Dataset with PubFig. Performance of GM Attack ($\alpha$ = 0.1, $\epsilon$ = 16) on ResNet50 is reported. ACC of the clean model (%) is reported in the last row.
Poison pair
20 classes
(5445 images)
30 classes
(6626 images)
40 classes
(7631 images)
50 classes
(8554 images)
$1-7$ 100 (0.0) 99.7 (0.5) 100.0 (0.0) 98.0 (2.7)
$3-6$ 100 (0.0) 100 (0.0) 100.0 (0.0) 100.0 (0.0)
$5-1$ 99.1 (1.6) 93.8 (1.1) 96.0 (1.4) 93.5 (4.9)
$6-8$ 92.8 (3.8) 95.0 (4.1) 85.1 (6.6) 84.0 (9.9)
$7-3$ 99.7 (0.5) 98.5 (1.9) 100.0 (0.0) 92.5 (5.9)
ACC (%) 99.7 (0.1) 99.7 (0.2) 99.3 (0.1) 97.1 (0.2)
Table 14: Merged Dataset with CelebA. Attack success rate (%) of GM Attack ($\alpha$ = 0.1, $\epsilon$ = 16) on ResNet50 is reported. Prediction accuracy (%) is reported in the last row.
Poison pair
100 classes
(6227 images)
200 classes
(9227 images)
400 classes
(15227 images)
$1-6$ 97.4 (1.7) 95.7 (1.7) 96.6 (3.0)
$5-1$ 80.9 (6.7) 81.8 (6.2) 84.6 (7.9)
$6-8$ 93.1 (4.1) 86.2 (6.6) 51.5 (6.7)
$7-3$ 93.2 (4.8) 67.9 (10.9) 86.7 (2.1)
$9-5$ 93.1 (3.6) 94.2 (2.6) 59.1 (9.3)
Prediction accuracy (%) 96.4 (0.4) 93.6 (0.2) 93.4 (0.3)
Since the poison ratio $\alpha$ is defined over the target class dataset, the number of poisoned images does not change when we increase the dataset size. Interestingly, as shown in Table 14 and Table 13, scaling dataset size only minimally impacts the attack performance.
C.4 Evaluating Digital Trigger for CLPBA
In CLPBA, we assume that the attackers hold trigger samples from the source class to perform poisoning. However, this may not be practical in real-world scenarios when the attacker can access clean images from the source class but not the trigger images. To mitigate this, the attacker can inject a digital image of the physical trigger into the source-class images as an alternative. In this section, we aim to evaluate the performance of CLPBA when the poisoned samples are crafted with photoshopped trigger images. As in Table 15, using photoshopped images has varying effectiveness for different poison pairs.
Table 15: Evaluating the use of digitally-modified trigger images for GM attack. The metric reported is ASR (%)
Poison pairs 1-7 2-0 6-8 9-5
GM-digital 98.0 (1.0) 0.0 (0.0) 68.6 (8.6) 66.3 (21.5)
GM-physical 100.0 (0.0) 99.4 (1.0) 98.3 (0.8) 97.8 (0.2)
C.5 Evaluating transferability of CLPBA
We evaluate the black-box settings of CLPBA for two poison pairs $7-1$ and $3-6$ in Table 16 and Table 17 Generally, we find that CLPBA under black-box settings have a worse performance, which means that the semantic features extracted by different model architectures may differ greatly. The same attack on different poison pairs may have vastly different results. For example, the poison pair $3-6$ with sunglasses trigger has a much better transferability than the poison pair $7-1$. This shows that the effectiveness of the trigger depends on the alignment between the trigger and natural features of the target class in the physical space.
Table 16: Evaluate transferability of CLPBA on poison pair $1-7$
Attacker Victim Inception-ResNetV1 ResNet50 VGG16
GM FM GM FM GM FM
Inception-ResNetV1 96.9 0.3 46.4 16.2 20.8 5.4
ResNet50 1.7 0.3 99.7 92.0 35.3 6.3
VGG16 3.7 0.3 91.5 24.2 100.0 83.2
Table 17: Evaluate transferability of CLPBA on poison pair $3-6$
Attacker Victim Inception-ResNetV1 ResNet50 VGG16
GM FM GM FM GM FM
Inception-ResNetV1 100.0 79.2 30.4 18.8 99.5 97.6
ResNet50 72.9 55.6 100.0 92.3 98.6 98.1
VGG16 72.9 49.3 71.5 22.2 100.0 100.0
Appendix D Additional Visualization
Figure 13: Additional visualization of poisoned samples
Visualization of poisoned samples. As shown in Figure 13, the perturbations of GM appear to reconstruct the sunglasses trigger. In FM, the perturbations are more uniform and fine-grained, but in some poisoned samples the perturbations resemble those crafted by GM (the image in the last column). We believe that the reconstruction of the sunglasses trigger is the key to backdoor the victim model. However, not all samples exhibit the perturbation pattern of a trigger. GM is more successful than FM because it successfully reconstructs the trigger in the perturbation in most of the samples.
Figure 14: Additional visualizations of GM attack with $\ell_{\infty}$ regularization and Pixel Regularization. Left image generated by standard $\ell_{\infty}$ regularization, Right Image generated by Pixel Regularization.
Figure 15: Visualization of poisoned squirrel images with the standard $\ell_{\infty}$ regularization
Figure 16: Visualization of poisoned squirrel images with pixel regularization
Appendix E Defenses
To defend against backdoor attacks in DNNs, a multitude of strategies have been developed, each addressing different vulnerabilities. Preprocessing-based defenses [30, 10] aim to weaken embedded triggers by altering the training data, while Model reconstruction-based defenses [29, 59] rectify the poisoned model by eliminating hidden backdoors through retraining, neuron pruning, or knowledge distillation. Trigger synthesis-based defenses [51, 7] counteract backdoors by identifying and neutralizing the specific triggers used in attacks. Model diagnosis-based defenses [22, 55] scrutinize a model’s behavior to detect anomalies indicative of a backdoor. Filtering-based defenses [49, 3, 14, 37, 17] identify and filter suspicious data that are likely to be poisoned by adversaries. Firewall-based defenses aim to detect malicious inputs at inference time and discard them so that the trained backdoored model will not behave maliciously.
In the paper, we evaluate CLPBA against four representative filtering-based defenses and four representative firewall defenses:
  • Activation Clustering (AC) [3]: This defense filters poisoned inputs in the latent space of the poisoned model via clustering. It assumes that the poisoned inputs form a small cluster separate from the clean inputs.
  • Spectral Signatures (SS) [49]: This defense identifies a common property, spectral signature, of backdoor attacks: Feature representations of the poisoned samples strongly correlate with the top singular vector of the feature covariance matrix. This defense then filters a predefined number of samples that have the highest correlation to the singular vector.
  • SPECTRE [17]: This defense improves upon the Spectral Signature defense with robust covariance estimation that amplifies the spectral signature of corrupted data.
  • DeepKNN [37]: This defense was originally introduced for clean-label data poisoning. It assumes that poisoned samples exhibit different feature distributions from clean examples in the feature space. It then uses K-nearest neighbors to filter samples with the most number of conflicting neighbors (neighbors that have different labels).
  • STRIP [14]: This defense detects poisoned inputs by applying random perturbations and observing the model’s prediction entropy. Poisoned inputs typically produce more consistent (lower entropy) predictions under perturbations compared to clean inputs.
  • IBD-PSC [19]: This defense clusters inputs based on their feature representations and identifies poisoned samples as distinct clusters in the feature space, separate from clean samples.
  • Frequency-based Detection [56]: This defense identifies backdoor triggers by analyzing frequency patterns in the input data. Trigger artifacts often show statistically distinct patterns that can be isolated through frequency domain analysis.
  • Cognitive Distillation [20]: This method detects backdoor patterns by isolating minimal features, called Cognitive Patterns (CPs), that trigger the same model output. Backdoor samples consistently yield unusually small CPs, making them easy to identify.
E.1 Evaluation on filtering defenses.
Backdoor poisoning attacks have been observed to leave discernible traces in the feature space of poisoned models. [49, 17, 3, 38]. By projecting feature representations of target samples (containing both clean and poisoned samples) on the first two principle axes (Figure 17), we can see a clear distinction between the poisoned samples and clean samples. The separation exists because the perturbations on poisoned samples contain unnatural trigger features of $\widetilde{D}_{s}$, from which the model can learn as features of the target class. However, from Table 6 in the main paper, we can still observe that the GM attack is still effective despite the filtering defenses having high TPR. This shows that only a few effective instances are meaningful for the attack.
Figure 17: 2D-PCA view of target-class representation produced by the poisoned model
E.2 Evaluation on firewall defenses.
As demonstrated in Table 6 in the main paper, input-level detection methods are not effective for CLPBAs because they either miss trigger samples or incorrectly filter out too many benign samples. This behavior is expected, as CLPBAs challenge the main assumptions underlying these defenses:
  • STRIP and IBD-PSC: These methods assume that backdoor correlation (trigger and target label prediction) is more consistent than the classification of benign samples, and thus find ways to unlearn normal classification tasks to highlight trigger samples. However, since CLPBAs work by synthesizing natural features from the target class with the trigger, unlearning normal classification tasks also unlearns the backdoor correlation between the trigger and the target class.
  • Frequency-based Detection: This method assumes that poisoned samples exhibit high-frequency artifacts that differ from benign ones. This assumption holds true for digital triggers, where there is no inherent correlation between the trigger and the natural image content in the pixel space. However, physical triggers, which are integrated naturally into the image, do not produce such high-frequency artifacts, making this defense less effective.
  • Cognitive Distillation: This approach assumes that a backdoored model focuses on much smaller regions for classifying trigger samples than for classifying clean samples. However, because CLPBAs aim to embed the distribution of the source class with the trigger into the feature space of the target class, the classification region for trigger samples is larger. The model relies on a combination of the trigger and natural features of the source class for misclassification, as evidenced by the Grad-CAM images in Figure 4 (Appendix A).
It is also important to note that these defenses are built for dirty-label all-to-one attacks instead of clean-label one-to-one attacks as CLPBA. Therefore, the assumption of a consistent backdoor correlation for STRIP and IBD-PSC may not hold true.
Appendix F Limitations and future works
High variance. Despite using identical training data and model structures, ASR can fluctuate dramatically, peaking in one training session and then significantly declining in another with a different seed. Such high variance is a common challenge in clean-label backdoor attacks due to the randomness of model training and the complexities of optimizing the attacker’s poisoning objectives, such as gradient alignment (Equation 9). This underscores the challenge of maintaining consistent attack efficacy in real-world applications.
Dataset. Our study is limited by the lack of diversity in the collected human facial dataset. The dataset includes only 10 Asian people, 9 males & 1 female, all within the 20-25 year age range because we are unable to find suitable volunteers with diversity in terms of age, race, and gender. While we aimed to collect a dataset that reflects real-world scenarios by incorporating physical randomness in the photos (e.g., different lighting conditions, shooting angles, cameras, etc.), our dataset may still not be diverse enough to cover real-world settings. The lack of diversity in race, gender, age demographics, and physical variability might impact the generalizability of our findings to real-world scenarios.
Future work. Our study focuses solely on Image Classification tasks. A natural direction from this work would be to study clean-label physical backdoor attacks for different tasks, such as Object Detection and Image Segmentation.
Appendix G Societal impact
The challenges highlighted by the Clean-label Physical Backdoor Attack (CLPBA) can push the AI community towards developing more trustworthy AI models. By identifying and demonstrating the vulnerabilities associated with CLPBA, this research can drive the development of more advanced and effective defense mechanisms against CLPBA. The findings can be used to educate developers, security professionals, and policymakers about the risks of CLPBA and the importance of securing AI systems. Increased awareness can lead to better training programs and a more security-conscious development culture. Security experts and researchers can use these insights to create robust detection and mitigation strategies, enhancing the overall security of AI systems. Also, policymakers can draft more effective regulations to safeguard against such threats.
Figure 18: Example images of the 10 volunteers, representing the 10 classes in the dataset.
Figure 19: Physical triggers of the animal classification dataset
Figure 20: Left column: Main trigger. Right columns: Three false triggers with matching shapes, sizes, colors, and locations on the face to test the false positive rate of the model.